OpenStudy (anonymous):

Verify that this is an Inverse. I know this will come out to be x but am a bit unsure where to go from here

OpenStudy (anonymous):

ok, it won't let me type an equation, so one minute

OpenStudy (anonymous):

http://gyazo.com/9781886559bfca347c8b4dc5f756b5b4

OpenStudy (phi):

prove it's the inverse of what ?

OpenStudy (anonymous):

my bad, I had already put the equations into each other. The two equations are both the same (2x+1)/(3x-2)

OpenStudy (phi):

if you have f(x) and the inverse g(x) then g(f(x)) = x and f(g(x)) = x so if you plugged f(x) into g(x) and got x, you showed that g(x) is the inverse function

OpenStudy (anonymous):

yah. I know. But i plugged it in, and got this http://gyazo.com/40d985609f5ddf098f636f6cebb026b5

OpenStudy (anonymous):

so i distributed the 2 (numerator) and the 3(denomonator), but I don't know where to to get from here

OpenStudy (phi):

Are you sure about your equations? if you start with f(x) = (x+1)/(3x-2) and g(x) =(2x+1)/(3x-2) *slight different from f(x)* and plug f(x) into g(x) , I think it will work out to x

OpenStudy (anonymous):

oh, ok, that makes sense, thanks

OpenStudy (phi):

nope, that is not correct...but I think there is a typo in the question somewheres...

OpenStudy (anonymous):

OK, so for the sake of it assume the equation that i posted originally was correct, does that simplify down to x? Because i feel like it should only i am not sure how to simplify it

OpenStudy (phi):

how about f(x)= (x+1)/(3x-2) then the inverse will be g(x) = (2x+1)/(3x-1)

OpenStudy (phi):

Yes, your very first post simplifies to x first step: multiply top and bottom by (3x-2)

OpenStudy (anonymous):

Ok, there we go. I couldn't figure out how to get rid of the fractions within the fraction, it all worked out now. Thank you